OpenStudy (anonymous):

A number of discs, each of momentum M kg m/s are striking a wall at the rate of n discs per minute. The force associated with these discs, in newtons, would be

OpenStudy (anonymous):

momentum divided by time equals to force the force=\[(M \times n)\]\[\div60\]

OpenStudy (anonymous):

\[(M \times n)\div60\]

OpenStudy (jamesj):

Right. Impulse = integral of Force over time = change in momentum ==> Force = (change in impulse)/( time) = (change of momentum) / (time)

OpenStudy (anonymous):

oh wait it should be 2(Mxn)÷60 because the change in momentum should be 2M

OpenStudy (anonymous):

Assuming it is elastic collision